OpenStudy (anonymous):

What is the formula for finding the surface area of a rectangular prism? Select one of the options below as your answer: A. 2(l × l + w × w + h × h) B. 2(l × w + l × h + w × h) C. l × w + l × h + w × h D. 4(l × w + l × h + w × h)

OpenStudy (cwrw238):

the area is the sum of the areas of each face of the prism

OpenStudy (anonymous):

(a, b, and c are the lengths of the 3 sides) In words, the surface area of a rectangular prism is the area of the six rectangles that cover it. But we don't have to figure out all six because we know that the top and bottom are the same, the front and back are the same, and the left and right sides are the same. The area of the top and bottom (side lengths a and c) = a*c. Since there are two of them, you get 2ac. The front and back have side lengths of b and c. The area of one of them is b*c, and there are two of them, so the surface area of those two is 2bc. The left and right side have side lengths of a and b, so the surface area of one of them is a*b. Again, there are two of them, so their combined surface area is 2ab.

OpenStudy (cwrw238):

and there are 3 pairs of faces each pair ( eg w x l) have the same area

OpenStudy (anonymous):

The bottom side and the top side are equal and have l and w as dimensions The area for the top and bottom side is l× w + l × w = 2 × l × w The front side (shown in sky blue) and the back side (not shown) are equal and have h and l as dimensions The area for the front and the back side is l× h + l × h = 2 × l × h Then, the last two sides have h and w as its dimensions. One side is shown in purple The area for the front and the back side is w× h + w × h = 2 × w × h The total surface area, call it SA is: SA = 2 × l × w + 2 × l × h + 2 × w × h

OpenStudy (anonymous):

ohh thanks guys :)

OpenStudy (anonymous):

give me a medal